# An information-matching approach to optimal experimental design and active learning

### v3

@@ Abstract
However, many applications of predictive models do not require precise parameter estimates per se, but accurate predictions for key quantities of interest (QoIs) [18, 32, 33, 34, 35]. This distinction is well illustrated by sloppy models, where many parameter combinations are practically unidentifiable, yet still yield precise predictions [36, 37, 38]. Such models typically have only a few stiff, identifiable directions, while the rest are sloppy and poorly constrained. Furthermore, the parameters that are identifiable from the training data may not align with those most relevant for predicting QoIs. Precise predictions are still possible when these QoI-relevant directions fall within the identifiable subspace; however, if they do not, the resulting uncertainty may be large or even divergent, regardless of how well some parameters are constrained [39].
@@
Motivated by these considerations, we develop an information-matching method for OED that prioritizes the precision of predictions for the target QoIs. This approach uses a model parameterized by $\boldsymbol{\theta}$ in two key scenarios: training, $\bm{f}$, and prediction, $\bm{g}$ (see Fig. 1). Given a dataset of $M$ independent inputs $\{\mathbf{x}_{m}\}_{m=1}^{M}$ and their corresponding ground truth labels $\{\mathbf{p}_{m}\}_{m=1}^{M}$, we first use $\bm{f}(\boldsymbol{\theta};\mathbf{x}_{m})$ to train the parameters against the data $\{\mathbf{x}_{m},\mathbf{p}_{m}\}$. Then, we use the trained parameters to predict the target QoIs $\mathbf{q}$ corresponding to the input $\mathbf{y}$ through the mapping $\bm{g}(\boldsymbol{\theta};\mathbf{y})$. In this scenario, $\mathbf{y}$ may act as an input, similar to $\mathbf{x}_{m}$, as a control parameter for the QoIs, or as a discrete index to distinguish between different QoI values. The information-matching method leverages both scenarios to identify the minimal subset of training data that contains the information necessary to precisely constrain the parameters relevant to target QoIs. Our strategy is: given a target precision for target QoIs, align the FIM for the training data with that of the QoIs, ensuring the training data carry the information needed to constrain the predictions precisely. Thus, only the parameter combinations that need to be identified are trained, bypassing numerical stability issues in cases where the FIMs are ill-conditioned.
Figure 1: The information-matching workflow. The panels represent (middle) a model parameterized by $\boldsymbol{\theta}$, (left) the target QoIs to predict, and (right) the dataset for training. First, one selects the target precision for the QoIs (blue envelope in the left panel). This QoI precision induces a minimal confidence region in parameter space: blue ellipse in the center panel. The information matching criterion selects training data and target precision (right panel) such that the resulting parameter uncertainty (yellow ellipse in the center panel) is more restrictive than that induced by the QoIs. Propagating the parameter uncertainty to the QoIs gives predictions that are at least as precise as the original target. This relationship holds even if the target uncertainties become infinite for certain inputs (dashed blue curves in the left panel), resulting in the target parameter confidence diverging for some parameter combinations (dashed blue ellipse in the middle panel, extending in some directions).
@@
The information-matching method leverages the FIMs of both the training data and target QoIs (Eqs. (2) and (3), respectively) to guarantee that the information in the training data is sufficient to achieve the target precision in the QoIs. To select the minimal set of training data that achieves this minimal information bound, we solve the following convex problem for the weight vector $\mathbf{w}=\begin{bmatrix}w_{1}&w_{2}&\dots&w_{M}\end{bmatrix}^{T}$:
@@
The matrix inequality constraint in Eq. (4) is crucial for ensuring the target precision on the QoIs. Formally, it means that the difference $\mathcal{I}-\mathcal{J}$ is positive semidefinite. Intuitively, it indicates that fitting the down-selected data results in smaller parameter variance compared to fitting the target QoIs directly, as illustrated in Fig. 1. Theorem 1 formalizes this statement (proof in the SM).
@@
Let $\bm{g}(\boldsymbol{\theta};\mathbf{y})$ denote a mapping from the model parameters $\boldsymbol{\theta}$ to the target QoIs for input $\mathbf{y}$ that is analytic at $\boldsymbol{\theta}_{0}=\left<\boldsymbol{\theta}\right>_{\boldsymbol{\theta}}$, where $\left<\cdot\right>_{\boldsymbol{\theta}}$ denotes an expectation value over the distribution of parameters. Consider parameters of the form $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}+\epsilon\delta\boldsymbol{\theta}$. If the constraints in Eq. (4) are satisfied, then
@@
We use the IEEE 39-bus system [40], represented graphically in Fig. 2, as a benchmark. Nodes denote buses, and edges represent transmission lines and transformers [3]. The parameters $\boldsymbol{\theta}$ to be inferred are the system states (voltage magnitude and angle), and $\bm{f}$ represents PMU measurements. Full parameter identifiability implies that the target QoIs are the state variables themselves $\bm{g}(\boldsymbol{\theta};\mathbf{y})=\boldsymbol{\theta}$] and the target FIM is non-singular. We enforce this condition by setting $\mathcal{J}=\lambda I$ for some small $\lambda>0$. The information-matching condition implies that $\mathcal{I}$ is also non-singular, and all states are observable. The optimal PMU placements for this problem are represented as the orange highlighted buses in Fig. 2. Our results on this initial benchmark test align with previous studies [4, 5, 3]. Even without preassigning PMU locations, this method naturally selects the same buses identified elsewhere [6, 7].
Figure 2: The IEEE 39-bus system. Heavy black lines denote buses (nodes), green lines with arrows indicating power flow denote transmission lines (edges), w-shaped symbols represent transformers, and circles are generators. Orange buses are optimal PMU locations for complete network observability. Other colors (red, green, or blue) indicate the optimal placements for partial observability in the corresponding area. Many buses are double-highlighted with orange and another color, showing overlaps between full and partial observability. Non-overlapping optimal buses result from unobserved branches.
@@
To localize the two sources using Eq. (4), the target FIM $\mathcal{J}$ is a diagonal matrix, where diagonal elements that correspond to source locations are set to their inverse target precision, and all other elements are zero. The optimal receiver locations are shown as the large dots in Fig. 3, which account for only 5% of the candidate locations.
@@
We apply Algorithm 1 to develop an optimal 15-parameter Stillinger–Weber (SW) potential for molybdenum disulfide (MoS2) to precisely predict the energy ($E$) as a function of lattice parameter ($a$) under uniform in-plane strain [33, 39, 49, 50]. The predictions are shifted by $E_{c}$ (the energy at the equilibrium lattice constant $a_{0}$) to align the minimum with the origin, effectively showing strain-induced energy changes. The candidate dataset comprises 2000 atomic configuration snapshots from an ab-initio molecular dynamics trajectory at 750 K, each with 96 Mo and 192 S atoms [33, 51]. While this dataset contains force labels, in practice, the candidate dataset does not need to include labels; Algorithm 1 generates them on demand for the optimal configurations. We choose the target precision to be 10% of the values predicted by the potential trained on the full dataset [33, 51]. In Fig. 4, we compare the uncertainties of the target QoI obtained from the optimal configurations (red envelope) with the target uncertainty (blue envelope). Our findings indicate that seven atomic configurations suffice to constrain the parameters and achieve the target precision. Additional results for other SW potentials for silicon are also provided in the SM.
@@
We have demonstrated the versatility of the approach across diverse domains, including power system networks, underwater acoustics, and AL for interatomic potential development. We have also explored the hypothesis that relatively few data points may efficiently constrain typical QoIs. The potential applications of this method extend beyond these areas, encompassing domains such as biology, neuroscience, geology, and atmospheric science. Future work may extend the method to larger models and machine-learning applications, including machine-learned interatomic potentials. Investigating the feasibility and potential advantages of such integration could unlock broader applications and insights.

### v4

@@ Abstract
However, many applications of predictive models do not require precise parameter estimates per se, but accurate predictions for key quantities of interest (QoIs) [10, 42, 47, 35, 39]. This distinction is well illustrated by sloppy models, where many parameter combinations are practically unidentifiable, yet still yield precise predictions [44, 43, 7, 33]. Such models typically have only a few stiff, identifiable directions, while the rest are sloppy and poorly constrained. Furthermore, the parameters that are identifiable from the training data may not align with those most relevant for predicting QoIs. Precise predictions are still possible when these QoI-relevant directions fall within the identifiable subspace; however, if they do not, the resulting uncertainty may be large or even divergent, regardless of how well some parameters are constrained [24].
@@
Figure 1: Relationship between training data (left), model parameters (middle), and QoIs (right) in the information-matching framework. One first selects the target precision for the QoIs (blue envelope in the right panel). This QoI precision induces a minimal confidence region in parameter space (blue ellipse in the middle panel). The information-matching criterion selects training data and target precision (orange envelope in the left panel) such that the resulting parameter uncertainty (orange ellipse in the middle panel) is more restrictive than that induced by the QoIs. Propagating the parameter uncertainty to the QoIs gives predictions that are at least as precise as the original target (orange envelope in the right panel). This relationship holds even if the target uncertainties were divergent for certain QoIs (dashed blue curves in the right panel), resulting in the target parameter confidence diverging for some parameter combinations (dashed blue ellipse in the middle panel, extending in some directions).
Motivated by these considerations, we develop an information-matching method for OED that prioritizes the precision of predictions for the QoIs. This approach uses a model parameterized by $\boldsymbol{\theta}$ in two key scenarios: training, $\bm{f}$, and prediction, $\bm{g}$ (see Fig. 1). Given a dataset of $M$ independent inputs $\{\mathbf{x}_{m}\}_{m=1}^{M}$ and their corresponding ground truth labels $\{\mathbf{p}_{m}\}_{m=1}^{M}$, we first use $\bm{f}(\boldsymbol{\theta};\mathbf{x}_{m})$ to train the parameters against the data $\{\mathbf{x}_{m},\mathbf{p}_{m}\}$. Then, we use the trained parameters to predict the QoIs $\mathbf{q}$ corresponding to the input $\mathbf{y}$ through the mapping $\bm{g}(\boldsymbol{\theta};\mathbf{y})$. In this scenario, $\mathbf{y}$ may act as an input, similar to $\mathbf{x}_{m}$, as a control parameter for the QoIs, or as a discrete index to distinguish between different QoI values. The information-matching method leverages both scenarios to identify the minimal subset of training data that contains the information necessary to precisely constrain the parameters relevant to QoIs. Our strategy is: given a target precision for QoIs, align the FIM for the training data with that of the QoIs, ensuring the training data carry the information needed to constrain the predictions precisely. Thus, only the parameter combinations that need to be identified are trained, bypassing numerical stability issues in cases where the FIMs are ill-conditioned.
@@
The information-matching method leverages the FIMs of both the training data and QoIs (Eqs. (2) and (3), respectively) to guarantee that the information in the training data is sufficient to achieve the target precision in the QoIs. To select the minimal set of training data that achieves this minimal information bound, we solve the following convex problem for the weight vector $\mathbf{w}=\begin{bmatrix}w_{1}&w_{2}&\dots&w_{M}\end{bmatrix}^{T}$:
@@
The matrix inequality constraint in Eq. (4) is crucial for ensuring the target precision on the QoIs. Formally, it means that the difference $\mathcal{I}-\mathcal{J}$ is positive semidefinite. Intuitively, it indicates that fitting the down-selected data results in smaller parameter variance compared to fitting the QoIs directly, as illustrated in Fig. 1. Theorem 1 formalizes this statement (proof in the SM).
@@
Let $\bm{g}(\boldsymbol{\theta};\mathbf{y})$ denote a mapping from the model parameters $\boldsymbol{\theta}$ to the QoIs for input $\mathbf{y}$ that is analytic at $\boldsymbol{\theta}_{0}=\left<\boldsymbol{\theta}\right>_{\boldsymbol{\theta}}$, where $\left<\cdot\right>_{\boldsymbol{\theta}}$ denotes an expectation value over the distribution of parameters. Consider parameters of the form $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}+\epsilon\delta\boldsymbol{\theta}$. If the constraints in Eq. (4) are satisfied, then
@@
Figure 2: The IEEE 39-bus system. Buses are represented by thick black lines, while transmission lines and transformers are shown as thin lines connecting them. Loads are depicted as black arrows pointing outward from the buses, and the circled labels G1 through G10 indicate the generators. Buses highlighted in orange denote the optimal PMU placements for full observability of the entire network. Buses highlighted in other colors (red, green, and blue) represent the optimal PMU placements for partial observability in the corresponding area. Many buses are double-highlighted with orange and another color, showing overlaps between full and partial observability. Non-overlapping optimal buses result from unobserved branches.
We use the IEEE 39-bus system [18], represented graphically in Fig. 2, as a benchmark. Nodes denote buses, and edges represent transmission lines and transformers [53]. The parameters $\boldsymbol{\theta}$ to be inferred are the system states (voltage magnitude and angle), and $\bm{f}$ represents PMU measurements. Full parameter identifiability implies that the QoIs are the state variables themselves $\bm{g}(\boldsymbol{\theta};\mathbf{y})=\boldsymbol{\theta}$] and the QoI FIM is non-singular. We enforce this condition by setting $\mathcal{J}=\lambda I$ for some small $\lambda>0$. The information-matching condition implies that $\mathcal{I}$ is also non-singular, and all states are observable. The optimal PMU placements for this problem are represented as the orange highlighted buses in Fig. 2. Our results on this initial benchmark test align with previous studies [30, 2, 53]. Even without preassigning PMU locations, this method naturally selects the same buses identified elsewhere [28, 11].
@@
To localize the two sources using Eq. (4), the QoI FIM $\mathcal{J}$ is a diagonal matrix, where diagonal elements that correspond to source locations are set to their inverse target precision, and all other elements are zero. The optimal receiver locations are shown as the large dots in Fig. 3, which account for only 5% of the candidate locations.
@@
We apply Algorithm 1 to develop an optimal 15-parameter Stillinger–Weber (SW) potential for molybdenum disulfide (MoS2) to precisely predict the energy ($E$) as a function of lattice parameter ($a$) under uniform in-plane strain [47, 24, 49, 48]. The predictions are shifted by $E_{c}$ (the energy at the equilibrium lattice constant $a_{0}$) to align the minimum with the origin, effectively showing strain-induced energy changes. The candidate dataset comprises 2000 atomic configuration snapshots from an ab-initio molecular dynamics trajectory at 750 K, each with 96 Mo and 192 S atoms [47, 50]. While this dataset contains force labels, in practice, the candidate dataset does not need to include labels; Algorithm 1 generates them on demand for the optimal configurations. We choose the target precision to be 10% of the values predicted by the potential trained on the full dataset [47, 50]. In Fig. 4, we compare the uncertainties of the QoI obtained from the optimal configurations (red envelope) with the target uncertainty (blue envelope). Our findings indicate that seven atomic configurations suffice to constrain the parameters and achieve the target precision.
While some details of the calculation—the selected configurations, optimal weights, and optimal parameters—depend on the choice of initial parameters, these differences do not affect the validity of the final estimated uncertainty. As long as the problem is feasible, the resulting prediction uncertainty estimated by the FIM is guaranteed to be within the target uncertainty. This dependence on initialization is examined in more detail in the SM, where we provide numerical examples based on the SW MoS2 model. Additional results for other SW potentials for silicon are also provided in the SM.
@@
We have demonstrated the versatility of the approach across diverse domains, including power system networks, underwater acoustics, and AL for interatomic potential development. We have also explored the hypothesis that relatively few data points may efficiently constrain typical QoIs. The potential applications of this method extend beyond these areas, encompassing domains such as biology, neuroscience, geology, and atmospheric science.
Despite these advantages, successfully solving Eq. (4) depends on the richness of the initial candidate dataset; if any parameter that is relevant for the QoIs cannot be constrained by any of the candidate datasets, the optimization is infeasible. Because Eq. (4) is convex, algorithms will detect when a problem is infeasible. In such a case, it would be desirable to augment the candidate datasets with more informative candidates. Identifying such additional data points to achieve feasibility remains an open, problem-dependent question. Our current practical approach is to generate large and diverse candidate datasets. This can often be done quickly and cheaply. For example, in interatomic potential development, many atomic configurations with varied lattice parameters can be generated, while in sensor placement problems, one may define a fine grid of potential sensor locations. Because Eq. (4) is convex and scales favorably with problem size, this approach is practical in many scenarios.
Future work may extend the method to larger models and machine-learning applications, including machine-learned interatomic potentials. Investigating the feasibility and potential advantages of such integration could unlock broader applications and insights. Finally, a deeper theoretical analysis of the limiting behavior of the optimization may offer valuable guidance on robustness and further broaden applicability across diverse scientific domains.